OpenStudy (anonymous):

Indentify the open interval(s) on which the function is increasing or decreasing: F(X)=(x^2)ln(x/2)

OpenStudy (mathmale):

1) find the first derivative of the given function. 2) set this derivative = to 0 and solve the resulting equation for x. These are your "critical values." 3) draw a horizontal line (your x-axis" and plot the zeros ("critical values"). Example:

OpenStudy (mathmale):

Here, with two critical values, we have 3 intervals. Choose 1 representative x-value from within each of the 3 intervals. Determine the sign of the 1st derivative at each of these representative values. Your results? Please share a graph (use the Draw utility, below).

OpenStudy (anonymous):

The derivative is f'(X)=2ln(X/2)X+X

OpenStudy (anonymous):

I can't set the equation to zero?

OpenStudy (anonymous):

0=2ln(X/2)X+X

OpenStudy (mathmale):

Buddy: mind taking a second look at the given function and trying again to find its derivative? \[f(x)=x^2*\ln \frac{ x }{ 2 }\]

OpenStudy (mathmale):

Have you already recognized that this function is a product? ...or that ln (x/2) can be expanded using the rule of logs for quotients?

OpenStudy (anonymous):

Let me try again

OpenStudy (anonymous):

It is (log(X))/2

OpenStudy (mathmale):

Rather than answer Yes or No: ln (x/2) = ln x - ln 2 and if \[f(x)=x^2*\ln \frac{ x }{ 2 }\]

OpenStudy (mathmale):

then the derivative is \[f'(x)=x^2*\frac{ d }{ dx }[\ln x-\ln 2]+[lnx-\ln2]*2x\]

OpenStudy (anonymous):

I got the derivative is 2xlb(X)+X

OpenStudy (mathmale):

Here I've used the product rule and the identity for the ln of a quotient. Pls go thru this and ask any questions necessary to clarify this process for yourself.

OpenStudy (anonymous):

I mean 2xln(X)+X

OpenStudy (anonymous):

Do I set that to zero now?

OpenStudy (mathmale):

My result is slightly different. We have to reconcile our two results.\[f'(x)=x+(2x)(lnx-\ln2)\]

OpenStudy (mathmale):

I'm letting you decide whether our results are equivalent or not. If they are, then yes, you set the first deriv. = to 0.

OpenStudy (anonymous):

I am having trouble setting it to 0 again

OpenStudy (anonymous):

2x(ln(X)-ln(2))+X=0

OpenStudy (anonymous):

I have X=0 and X=1.2131

OpenStudy (anonymous):

:(